OpenStudy (anonymous):

find the x-intercepts of the parabola with vertex (1,-108) and y-intercept(0,-105). PLEASE write anwser in the form of (x1,y1)(x2,y2) if it is necessary you can round to the nearest hundreth. <3 PLEASE HELP ME URGENT !

OpenStudy (anonymous):

y+108 = 3(x-1)^2 gives us ...(7,0) and (-5,0)

OpenStudy (anonymous):

y+108 = 3(x-1)^2 gives us ...(7,0) and (-5,0)

OpenStudy (anonymous):

are you sure love ? this is a very important test .

OpenStudy (anonymous):

seems fine ...do a rough check yourself ...note do not use this to cheat in a test! this is for reference only

OpenStudy (anonymous):

it was wrong :( but thanks ; 3 people thought the same thing . its ok :)